OpenStudy (anonymous):

If Jacob wrote a 5-page paper in 4 hours and Olivia wrote a 6-page paper in 5 hours, who wrote at a faster rate? Show your work.

OpenStudy (anonymous):

You just have to compare the fractions 5/4 and 6/5. Begin with getting a common denominator.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

5/4=11/1 6/5=19/1

OpenStudy (anonymous):

No, you've got to multiply both the upper part and lower part by the same number for each fraction so that the lower part on both the fractions is the same: \[\frac{5}{4}\]=\[\frac{25}{20}\] and \[\frac{6}{5}\]=\[\frac{24}{20}\]

OpenStudy (anonymous):

so obviously 5/4 is greater than 6/5

OpenStudy (anonymous):

lol